Reports 1-1 of 1 Clear search Modify search
ayaka.shoda - 10:38 Friday 20 September 2019 (10635) Print this report
PEM further decoupled

After finding coherence remained at some resonsnt frequences with the simple decoupling method, I tried to decouple the YAW from LEN and PIT.

The result is here. (The other one attached is coherence between YAW and LEN/PIT.)

We do not observe the peaks at 0.4-0.5 Hz (Whole chain mode), 0.57 Hz (IM L), 0.65 Hz (TM L) in YAW cool
I think it is good that we can also remove the 0.4-0.5 Hz peaks from YAW signal without using the data at those peaks. (I used that values only at 0.15 Hz(YAW), 0.57 Hz, and 0.65 Hz.)

My code for this calculation is in /kagra/Dropbox/Personal/Shoda/PRM/L2A/OL_diag.ipynb, it is a spagetti code though. 
I will implement the new matrix when the suspension available.

Images attached to this report
Comments to this report:
ayaka.shoda - 16:26 Tuesday 24 September 2019 (10680) Print this report

I updated the sensor input matrix while the PRM is misaligned.

Since the numbers are not changed drastically, I guess it does not affect the performance of the controls that much in the current situation.
Anyway we should check the spectra later when the PRM is aligned.

Search Help